OpenStudy (anonymous):

use the elimination method to solve the system of equations 5u+3v=-18 2u+v=-6

OpenStudy (stormfire1):

5u+3v=-18 2u+v=-6 Solve for one of the variables...I'll go with v v = -6-2u Now plug v back into the first equation: 5u+3(-6-2u)=-18 5u-18-6u=-18 5u-6u=0 -u = 0 Now plug u back into the second equation: 2(0)+v=-6 v = -6

OpenStudy (anonymous):

does anyone know how to factor x (x+3)-9(x+3)